OpenStudy (anonymous):

Simplify: (c-d)/(1/c - 1/d)

OpenStudy (anonymous):

This can be written in this form (see the attachment) and the answer is -cd

OpenStudy (saifoo.khan):

\[\Large \frac{c-d}{\frac{1}{c}-\frac{1}{d}}\]